OpenStudy (anonymous):

who is doing flvs geometry segment 2?

OpenStudy (anonymous):

im doing alg 2 on flvs.. Im better at geo than alg... are you good at alg? we can help each other

OpenStudy (anonymous):

sure i mean i could see what i can help you out with :)

OpenStudy (anonymous):

ok what is your question? and here is mine : Use graphing technology to find one of the approximate solutions of x2 = 5x + 8. Round to the nearest hundredth.

OpenStudy (anonymous):

Assuming that the 2 following the x term is an exponent, x^2 = 5x + 8 subtract 5x from both sides of the equal sign x^2 - 5x = 8 subtract 8 from both sides of the equal sign x^2 - 5x - 8 = 0 then use the quadratic formula to solve it and thats it

OpenStudy (anonymous):

mine is an activity haha so dont worry about it but thanks though

OpenStudy (anonymous):

ok... thank you though, goodluck!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

i am